OpenStudy (anonymous):

Without drawing the graph of the equation, determine how many points the parabola has in common with the x-axis and whether its vertex lies above, on, or below the x-axis. y = 3x2 – 12x + 12 How do I do this?

OpenStudy (anonymous):

if you factor out that expression, you'll see that the factor \(\large (x-2) \) has mulitplicity of 2 which means the graph will only touch the x-axis only at x=2.

OpenStudy (anonymous):

To determine how many points the parabola has in common find the discriminant. To find the discriminant use \[ b^2 - 4ac \] Now just plugin are values \[ -12^2 - 4(3)(12) = 0\] So our discriminant equals 0. Now that our discriminant is equal to 0 we can determine how man point the parabola has in common. If discriminant < 0 then there will be 0 points If discriminant > 0 then there will be 2 points If discriminant = 0 then there will be 1 point Now to fine the whether its vertex lies above, on, or below the x-axis we find the x coordinate of the vertex by using \[ x = -\frac{b}{2a}\] Now just plugin your number for the formula above. \[x = -\frac{-12}{2(3)} \] \[x = -\frac{-12}{2(3)} \] \[x =\frac{12}{6} \] \[x =2\] Now we know the x cord of our vertex so we can find y which will tell us if the vertex lies above, on, or below the x-axis. Just plugin 2 for x in the equation y = 3x2 – 12x + 12 \[ y = 3(2)^{2} – 12(2) + 12 \] \[ y = 12 – 24 + 12 \] \[ y = 0 \] Since y = 0 the vertex lies on the x-axis

OpenStudy (anonymous):

That means it's : 1 point in common; vertex on x-axis Right?